OpenStudy (anonymous):

in death valley, California the highest ground temp, recorded was 95'C on july 15, 1972. in the formula C=5/9(f-32), C represents the temp. in degrees Celsius and F represents the temp. in degrees Fahrenheit. to the nearest degree, what is the highest ground temp. in death valley in farhenheit?

OpenStudy (shamil98):

95 celsius??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i meant 94 C

OpenStudy (shamil98):

C=5/9(f-32) 94 = 5(f-32) / 9 846 = 5f - 160 1006 = 5f 201.2 farenheit